# Orbital Stability of Plane Waves in the Klein-Gordon Equation against Localized Perturbations

### v1

@@ 4. Energy estimates
$\begin{split}&2E(z_{1},z_{2})+C_{0}(\delta_{2})\lVert\theta_{\infty}^{\prime}\rVert_{L^{2}({\mathbb{R}})}^{2}+2|k+c\omega|\mathrm{e}^{2}\|\theta_{\infty}^{\prime}\|_{L^{1}({\mathbb{R}})}\\ &\qquad\geq(1-c^{2})\lVert z_{2}\rVert_{L^{2}({\mathbb{R}})}^{2}+\left(1-c^{2}\right)(1-C_{*}\delta_{2})\left(\lVert\rho^{\prime}\rVert_{L^{2}({\mathbb{R}})}^{2}+(1-\delta_{2})\lVert\theta^{\prime}\rVert_{L^{2}({\mathbb{R}})}^{2}\right)\\ &\qquad\qquad+\,\left(2a^{2}f^{\prime}(a^{2})-C_{*}\delta_{2}\right)\lVert\rho\rVert_{L^{2}({\mathbb{R}})}^{2}\\ &\qquad\qquad-\,\lvert k+c\omega\rvert\left(2+C_{*}\delta_{2}\right)\bigg{(}\gamma\lVert\rho\rVert_{L^{2}({\mathbb{R}})}^{2}+\frac{1}{\gamma}\lVert\theta^{\prime}\rVert_{L^{2}({\mathbb{R}})}\bigg{)},\end{split}$ (4.9)
where $C_{0}(\delta_{2})>0$ is a constant depending on $C_{*},c$ and $\delta_{2}$ only.
@@
First, we consider the case $f(a^{2})>0$. Here, (1.3) allows us to choose $c=-\frac{k}{\omega}\in(-1,1)$ such that $k+c\omega=0$. Subsequently choosing $\gamma=1$ and taking $\delta_{2}>0$ sufficiently small, we find that all constants in front of the norms on the right-hand side of (4.9) are strictly positive by (1.12). This settles the lower bound (4.4) in the case $f(a^{2})>0$.
@@
Subsequently taking $\delta_{2}>0$ sufficiently small, we again obtain that all constants in front of the norms on the right-hand side of (4.9) are strictly positive, yielding the bound (4.4) for the case $f(a^{2})<0$.
@@
Hence, taking $\delta_{2}>0$ sufficiently small and recalling (1.12), we find that all constants in front of the norms on the right-hand side of (4.9) are strictly positive. This establishes (4.4) for the case $f(a^{2})=0$, which concludes the proof. ∎
@@ 5. Proof of the main result
We prove our orbital stability result, Theorem 4. The proof relies on the conservation of the energy functional (4.1) and employs the upper and lower bounds obtained in Lemmas 9 and 11.
@@
Let $c\in(-1,1)$ be as in Lemma 11, and let $\delta>0$. Suppose $\theta_{\infty}\in{\mathcal{C}}^{1}({\mathbb{R}},{\mathbb{R}})$ is $a$-admissible and satisfies (1.14). Furthermore, let $w_{0}\in H^{2}({\mathbb{R}})$ and $v_{0}\in H^{1}({\mathbb{R}})$ satisfy (1.15).
@@
$z_{1}(y)=\frac{w_{0}(y)}{a\mathrm{e}^{\mathrm{i}ky-\mathrm{i}\theta_{\infty}(y)}},\qquad z_{2}(y)=\frac{1}{\sqrt{1-c^{2}}}\left(\frac{v_{0}(y)-\mathrm{i}\left(ck+\omega\right)w_{0}(y)+cw_{0}^{\prime}(y)}{a\mathrm{e}^{\mathrm{i}ky-\mathrm{i}\theta_{\infty}(y)}}-c\mathrm{i}\theta_{\infty}^{\prime}(y)\right).$
We note that $z_{1}\in H^{2}({\mathbb{R}})$ and $z_{2}\in H^{1}({\mathbb{R}})$. Moreover, there exists an $E_{0}$- and $E_{\infty}$-independent constant $C_{0}>0$ such that
$\|z_{1}\|_{H^{1}({\mathbb{R}})}\leq C_{0}E_{0}\leq C_{0}\delta,\qquad\|z_{2}\|_{L^{2}({\mathbb{R}})}\leq C_{0}(E_{0}+E_{\infty})\leq 2C_{0}\delta.$ (5.1)
@@
$u(x,t)=\left(1+z(x-ct,t)\right)a\mathrm{e}^{\mathrm{i}kx+\mathrm{i}\omega t-\mathrm{i}\theta_{\infty}(x-ct)}$ (5.2)
@@
Hence, an application of Proposition 8 yields an $E_{0}$- and $E_{\infty}$-independent constant $C_{1}>0$ and unique continuously differentiable functions $\rho,\theta\colon{\mathbb{R}}\times[0,t_{\max})\to{\mathbb{R}}$ with $\rho(0)=\rho_{0}$ and $\theta(0)=\theta_{0}$ such that the polar decomposition
@@
for all $t\in[0,t_{\max})$. The polar decompositions (5.3) and (5.5) along with the first bound in (5.6) allow us to apply Lemmas 9 and 11 to bound the right-hand side of (5.8) from above and the left-hand side of (5.8) from below, respectively. Recalling that the second coordinate of $Z(t)$ is given by $z_{t}(t)/\sqrt{1-c^{2}}$ and using (5.1), (5.4), and (5.7), this results in $E_{0}$- and $E_{\infty}$-independent constants $C_{2},C_{3}>0$ such that
$\begin{split}&\lVert z_{t}(t)\rVert_{L^{2}({\mathbb{R}})}^{2}+\lVert\rho(t)\rVert_{H^{1}({\mathbb{R}})}^{2}+\lVert\theta_{y}(t)\rVert_{L^{2}({\mathbb{R}})}^{2}\\ &\qquad\leq C_{2}\bigg{(}\lVert z_{2}\rVert_{L^{2}({\mathbb{R}})}^{2}+\lVert\rho_{0}\rVert_{H^{1}({\mathbb{R}})}^{2}+\lVert\theta_{0}^{\prime}\rVert_{L^{2}({\mathbb{R}})}^{2}+\lVert\theta_{\infty}^{\prime}\rVert_{L^{2}({\mathbb{R}})}^{2}+\lvert k+c\omega\rvert\,\lVert\theta_{\infty}^{\prime}\rVert_{L^{1}({\mathbb{R}})}\bigg{)}\\ &\qquad\leq C_{3}\left(E_{0}^{2}+E_{\infty}^{2}\right)\leq 2C_{3}\delta^{2}\end{split}$ (5.9)
@@
$\vartheta(x,t)=\theta(x-ct,t)+\theta_{\infty}(x)-\theta_{\infty}(x-ct)=\theta(x-ct,t)+\int_{-ct}^{0}\theta_{\infty}^{\prime}(x+y)\mathrm{d}y.$
@@
It remains to show (1.17). To this end, take $t\geq 0$ and $x_{*}\in{\mathbb{R}}$. We set $\psi(x,t):=\vartheta(x,t)-\theta_{\infty}(x)$, $\gamma=\psi(x_{*},t)$, and $I=[x_{*}-R,x_{*}+R]$. Using (5.9), (5.14), and
@@
we find an $E_{0}$- and $E_{\infty}$-independent constant $C_{5}>0$ such that
$\begin{split}\lVert\psi(t)-\gamma\rVert_{L^{2}(I)}&\leq\left(\int_{I}\int_{I}\int_{I}|\psi_{x}(y,t)||\psi_{x}(l,t)|\,\mathrm{d}y\mathrm{d}l\mathrm{d}x\right)^{\frac{1}{2}}\\ &\leq 2R\lVert\psi_{x}(t)\rVert_{L^{2}(I)}\leq C_{5}\left(E_{0}+E_{\infty}\right)\leq 2C_{5}\delta.\end{split}$ (5.15)
@@
$\left\lvert u_{x}(x,t)-a\mathrm{i}k\mathrm{e}^{\mathrm{i}kx+\mathrm{i}\omega t+\mathrm{i}\gamma}\right\rvert$ $\leq|k|\left\lvert u(x,t)-a\mathrm{e}^{\mathrm{i}kx+\mathrm{i}\omega t+\mathrm{i}\gamma}\right\rvert+C_{4}\left(\left\lvert\varrho_{x}(x,t)\right\rvert+\left\lvert\psi_{x}(x,t)\right\rvert\right)$
for $x\in I$. Taking $L^{2}$-norms of the latter two inequalities and applying (1.19) and (5.15) yields (1.17), thereby completing the proof. ∎
@@ 7. Spectral stability analysis
$\sigma_{c}=\left\{\lambda\in{\mathbb{C}}:\mathcal{P}_{c}(\lambda)\text{ is not boundedly invertible}\right\},$
@@
We say that the plane wave (1.2) is spectrally stable if $\sigma_{c}$ is confined to the imaginary axis and spectrally unstable if it is not spectrally stable. Since $\mathcal{P}_{c}(\lambda)$ has constant coefficients, we find that it is boundedly invertible if and only if its Fourier symbol
@@
is invertible for each $\ell\in{\mathbb{R}}$. That is, $\mathcal{P}_{c}(\lambda)$ is not boundedly invertible if and only if there exists $\ell\in{\mathbb{R}}$ such that $(\lambda,\ell)$ is a root of the determinantal function $\mathcal{E}_{c}\colon{\mathbb{C}}\times{\mathbb{R}}\to{\mathbb{C}}$ given by
@@
The following result, summarizing the outcome of our spectral analysis, shows that condition (1.12) is sharp in the sense that it coincides with a change in spectral stability. Its proof employs the instability index theory for quadratic operator pencils as developed in [BJK14].
@@
Assume that condition (1.12) holds. We show that the plane wave is spectrally stable. As in the orbital stability analysis, we choose $c\in{\mathbb{R}}$ depending on the sign of $f(a^{2})$. First, we consider the case $f(a^{2})>0$. Using (1.3) and setting $\smash{c=-\frac{k}{\omega}}\in(-1,1)$, we find that the diagonal self-adjoint operator $\mathcal{H}_{c}(\partial_{y})$ is positive semidefinite on $\smash{L^{2}_{\mathrm{per}}(0,L)}$ for any $L\geq 0$. Therefore, the instability index count in [BJK14] implies that the plane wave is spectrally stable. Next, we consider the case $\smash{f(a^{2})}<0$. We use (1.3) to set $\smash{c=-\frac{\omega}{k}}\in(-1,1)$. By (1.12) we find that
@@
for all $\ell\in{\mathbb{R}}$. Therefore, both eigenvalues of $\mathcal{H}_{c}(\mathrm{i}k\ell)$ are nonnegative and the self-adjoint operator $\mathcal{H}_{c}(\partial_{y})$ is positive semidefinite on $L^{2}_{\mathrm{per}}(0,L)$ for any $L\geq 0$. Therefore, the plane wave is spectrally stable by [BJK14].

### v2

@@ 4. Energy estimates
$\begin{split}2E(z_{1},z_{2})&\geq(1-c^{2})\lVert z_{2}\rVert_{L^{2}({\mathbb{R}})}^{2}+\left(1-c^{2}\right)(1-C_{*}\delta_{2})\left(\lVert\rho^{\prime}\rVert_{L^{2}({\mathbb{R}})}^{2}+\lVert\theta^{\prime}+\theta_{\infty}^{\prime}\rVert_{L^{2}({\mathbb{R}})}^{2}\right)\\ &\qquad+\,\left(2a^{2}f^{\prime}(a^{2})-C_{*}\delta_{2}\right)\lVert\rho\rVert_{L^{2}({\mathbb{R}})}^{2}\\ &\qquad-\,\lvert k+c\omega\rvert\left(2+C_{*}\delta_{2}\right)\bigg(\gamma\lVert\rho\rVert_{L^{2}({\mathbb{R}})}^{2}+\frac{1}{\gamma}\lVert\theta^{\prime}+\theta_{\infty}^{\prime}\rVert_{L^{2}({\mathbb{R}})}^{2}\bigg),\end{split}$ (4.9)
where $C_{0}(\delta_{2})>0$ is a constant depending on $C_{*},c$ and $\delta_{2}$ only. By Young’s inequality, we can estimate
$\lVert\theta^{\prime}+\theta_{\infty}^{\prime}\rVert_{L^{2}({\mathbb{R}})}^{2}\geq\frac{1}{2}\|\theta^{\prime}\|_{L^{2}({\mathbb{R}})}^{2}-3\|\theta_{\infty}^{\prime}\|_{L^{2}({\mathbb{R}})}^{2}.$ (4.10)
@@
First, we consider the case $f(a^{2})>0$. Here, (1.3) allows us to choose $c=-\frac{k}{\omega}\in(-1,1)$ such that $k+c\omega=0$. Subsequently choosing $\gamma=1$ and taking $\delta_{2}>0$ sufficiently small, we find that all constants in front of the norms on the right-hand side of (4.9) are strictly positive by (1.12). Combining the latter with (4.10), settles the lower bound (4.4) in the case $f(a^{2})>0$.
@@
Subsequently taking $\delta_{2}>0$ sufficiently small, we again obtain that all constants in front of the norms on the right-hand side of (4.9) are strictly positive, which, together with (4.10), yields the bound (4.4) for the case $f(a^{2})<0$.
@@
Hence, taking $\delta_{2}>0$ sufficiently small and recalling (1.12), we find that all constants in front of the norms on the right-hand side of (4.9) are strictly positive. Combining this with (4.10), establishes (4.4) for the case $f(a^{2})=0$, which concludes the proof. ∎
@@ 5. Proof of the main result
We prove our orbital stability result, Theorem 3. The proof relies on the conservation of the energy functional (4.1) and employs the upper and lower bounds obtained in Lemmas 7 and 9.
@@
Let $c\in(-1,1)$ be as in Lemma 9, and let $\delta>0$. Suppose $\theta_{\infty}\in{\mathcal{C}}^{1}({\mathbb{R}},{\mathbb{R}})$, $w_{0}\in H^{2}({\mathbb{R}})$, and $v_{0}\in H^{1}({\mathbb{R}})$ satisfy (1.14).
@@
$z_{1}(y)=\frac{w_{0}(y)}{a\mathrm{e}^{\mathrm{i}ky+\mathrm{i}\theta_{\infty}(y)}},\qquad z_{2}(y)=\frac{1}{\sqrt{1-c^{2}}}\left(\frac{v_{0}(y)-\mathrm{i}\left(ck+\omega\right)w_{0}(y)+cw_{0}^{\prime}(y)}{a\mathrm{e}^{\mathrm{i}ky+\mathrm{i}\theta_{\infty}(y)}}+c\mathrm{i}\theta_{\infty}^{\prime}(y)\right).$
We note that $z_{1}\in H^{2}({\mathbb{R}})$ and $z_{2}\in H^{1}({\mathbb{R}})$. Moreover, there exists an $E_{0}$-independent constant $C_{0}>0$ such that
$\|z_{1}\|_{H^{1}({\mathbb{R}})}\leq C_{0}E_{0}\leq C_{0}\delta,\qquad\|z_{2}\|_{L^{2}({\mathbb{R}})}\leq C_{0}E_{0}\leq 2C_{0}\delta.$ (5.1)
@@
$u(x,t)=\left(1+z(x-ct,t)\right)a\mathrm{e}^{\mathrm{i}kx+\mathrm{i}\omega t+\mathrm{i}\theta_{\infty}(x-ct)}$ (5.2)
@@
Hence, an application of Proposition 6 yields an $E_{0}$-independent constant $C_{1}>0$ and unique continuously differentiable functions $\rho,\theta\colon{\mathbb{R}}\times[0,t_{\max})\to{\mathbb{R}}$ with $\rho(0)=\rho_{0}$ and $\theta(0)=\theta_{0}$ such that the polar decomposition
@@
for all $t\in[0,t_{\max})$. The polar decompositions (5.3) and (5.5) along with the first bound in (5.6) allow us to apply Lemmas 7 and 9 to bound the right-hand side of (5.8) from above and the left-hand side of (5.8) from below, respectively. Recalling that the second coordinate of $Z(t)$ is given by $z_{t}(t)/\sqrt{1-c^{2}}$ and using (5.1), (5.4), and (5.7), this results in $E_{0}$-independent constants $C_{2},C_{3}>0$ such that
$\begin{split}&\lVert z_{t}(t)\rVert_{L^{2}({\mathbb{R}})}^{2}+\lVert\rho(t)\rVert_{H^{1}({\mathbb{R}})}^{2}+\lVert\theta_{y}(t)\rVert_{L^{2}({\mathbb{R}})}^{2}\\ &\qquad\leq C_{2}\bigg(\lVert z_{2}\rVert_{L^{2}({\mathbb{R}})}^{2}+\lVert\rho_{0}\rVert_{H^{1}({\mathbb{R}})}^{2}+\lVert\theta_{0}^{\prime}\rVert_{L^{2}({\mathbb{R}})}^{2}+\lVert\theta_{\infty}^{\prime}\rVert_{L^{2}({\mathbb{R}})}^{2}\bigg)\leq C_{3}E_{0}^{2}\leq 2C_{3}\delta^{2}\end{split}$ (5.9)
@@
$\vartheta(x,t)=\theta(x-ct,t)-\theta_{\infty}(x)+\theta_{\infty}(x-ct)=\theta(x-ct,t)-\int_{-ct}^{0}\theta_{\infty}^{\prime}(x+y)\mathrm{d}y.$
@@
It remains to show (1.16). To this end, take $R>0$, $t\geq 0$, and $x_{*}\in{\mathbb{R}}$. We set $\psi(x,t):=\vartheta(x,t)+\theta_{\infty}(x)$, $\gamma=\psi(x_{*},t)$, and $I=[x_{*}-R,x_{*}+R]$. Using (5.9), (5.14), and
@@
we find an $E_{0}$-, $t$-, $x_{*}$- and $R$-independent constant $C_{5}>0$ such that
$\begin{split}\lVert\psi(t)-\gamma\rVert_{L^{2}(I)}&\leq\left(\int_{I}\int_{I}\int_{I}|\psi_{x}(y,t)||\psi_{x}(l,t)|\,\mathrm{d}y\mathrm{d}l\mathrm{d}x\right)^{\frac{1}{2}}\\ &\leq 2\sqrt{R}\,\lVert\psi_{x}(t)\rVert_{L^{2}(I)}\leq C_{5}\sqrt{R}\,E_{0}\leq 2C_{5}\sqrt{R}\,\delta.\end{split}$ (5.15)
@@
[…]
for $x\in I$. Taking $L^{2}$-norms of the latter two inequalities and applying (1.18) and (5.15) yields (1.16), which completes the proof. ∎
@@ 7. Spectral stability analysis
$\sigma_{c}=\left\{\lambda\in{\mathbb{C}}:\mathcal{P}_{c}(\lambda)\text{ is not invertible}\right\},$
@@
We say that the plane wave (1.2) is spectrally stable if $\sigma_{c}$ is confined to the imaginary axis and spectrally unstable if it is not spectrally stable. Since $\mathcal{P}_{c}(\lambda)$ has constant coefficients, we find that it is invertible if and only if its Fourier symbol
@@
is invertible for each $\ell\in{\mathbb{R}}$. That is, $\mathcal{P}_{c}(\lambda)$ is not invertible if and only if there exists $\ell\in{\mathbb{R}}$ such that $(\lambda,\ell)$ is a root of the determinantal function $\mathcal{E}_{c}\colon{\mathbb{C}}\times{\mathbb{R}}\to{\mathbb{C}}$ given by
@@
The following result, summarizing the outcome of our spectral analysis, shows that condition (1.12) is sharp in the sense that it coincides with a change in spectral stability. Its proof employs the instability index theory for quadratic operator pencils as developed in [6].
@@
Assume that condition (1.12) holds. We show that the plane wave is spectrally stable. As in the orbital stability analysis, we choose $c\in{\mathbb{R}}$ depending on the sign of $f(a^{2})$. First, we consider the case $f(a^{2})>0$. Using (1.3) and setting $\smash{c=-\frac{k}{\omega}}\in(-1,1)$, we find that the diagonal self-adjoint operator $\mathcal{H}_{c}(\partial_{y})$ is positive semidefinite on $\smash{L^{2}_{\mathrm{per}}(0,L)}$ for any $L\geq 0$. Therefore, the instability index count in [6] implies that the plane wave is spectrally stable. Next, we consider the case $\smash{f(a^{2})}<0$. We use (1.3) to set $\smash{c=-\frac{\omega}{k}}\in(-1,1)$. By (1.12) we find that
@@
for all $\ell\in{\mathbb{R}}$. Therefore, both eigenvalues of $\mathcal{H}_{c}(\mathrm{i}k\ell)$ are nonnegative and the self-adjoint operator $\mathcal{H}_{c}(\partial_{y})$ is positive semidefinite on $L^{2}_{\mathrm{per}}(0,L)$ for any $L\geq 0$. Therefore, the plane wave is spectrally stable by [6].